OpenStudy (anonymous):

Can someone help me ! Delilah opens a credit card with an APR of 24.38% compounded monthly. How much is charged in interest this month if her balance is $550?

OpenStudy (tkhunny):

Have you considered dividing the annual rate by 12 to produce the monthly rate?

OpenStudy (anonymous):

24.38/12 = 2.031667% 2.031667% of 550 = 0.02031667*550 = 11.1741685 ?

OpenStudy (tkhunny):

Why do you doubt?